Find sinx/2 cosx/2 tanx/2 in sinx=1/4 x in second quardrant

Grade:11

1 Answers

Arun
6 years ago
Sinx = ¼, 
Hence cos x = \sqrt15/4
We know that cosx = 1- 2 sin²(x/2)
\sqrt15/4 = 1- 2 sin²(x/2)
Sin²(x/2) = (1/2) [1- \sqrt15/4]
Sin(x/2) = \LARGE \sqrt(1/2) [1- \sqrt15/4]
 
I think, now you will be able to solve other values